Please show every step

6) A police study tested eyewitness memory a week after a non-stressful interrogation of a cooperative suspect and a stressful interrogation of an uncooperative and belligerent suspect. The numbers of details recalled were recorded and the summary statistics are givern below: Groupn Non-stress 4053.3 11.6 Stress 4045.3 13.2 We want to construct a 95% confidence interval for the difference of mean amount of details recalled based on stressful or non-stressful interrogation methods. a) Are the samples independent or dependent? b) What are the conditions for constructing a 95% confidence interval? c) Construct the confidence interval d) State your conclusion in the context of the problem.
Answer #1

a)

Here we have two independent samples.

b)

Since sample sizes are greater 30 so conditions for constructing a 95% confidence interval has been satisfied.

c)

Here we have following information:

T1 = 53.3. n 1 = 40, sı = 1 1.6:T2 = 45.3. n2 = 40, s-= 13.2

The pooled standard deviation:

n1-1)s +(n2-1) s2 = 12.4258 rl

-------------

So standard error for difference in population mean is

52 + = 2.7785 S-1-T2

Degree of freedom for t is df=40+40-2=78, so t-critical value using excel function "=TINV(0.05,78)" is 1.991. So required confidence interval for mu_{1} - mu_{2} will be

(T1 - T2)ttcritcal S (8) t1.991 2.7785 8 t5.532

So required confidence interval is (2.468, 13.532).

d)

Since confidence interval does not contain zero so we can conclude that there is a significant difference between the population means.